OpenStudy (anonymous):

Unit 6 Rational Expression Multiply. (a - 4) x 2a a^2 - 4a (a + 3)(a - 1) A. 2a^2 a - 1 B. 2 a - 1 C. 2a a + 1 D. 2a 2 - 1

geerky42 (geerky42):

Can you take screenshot of problem? Because it's hard to read here...

OpenStudy (anonymous):

I can't it doesn't work I will trying and use equation

geerky42 (geerky42):

Why doesn't screenshot work? What computer are you using?/

OpenStudy (anonymous):

\[\frac{ (a - 4) }{ a^2 - 4a } x \frac{ 2a }{ (a + 3)(a - 1) }\]

geerky42 (geerky42):

You can factor \(a^2-4a\), just factor \(a\) out; \[a^2-4a\quad\Longrightarrow\quad a(a-4)\]

geerky42 (geerky42):

So you have \[\frac{ (a - 4) }{ a(a - 4) } \times \frac{ 2a }{ (a + 3)(a - 1) }\]

OpenStudy (anonymous):

okay so next is to cross out (a - 4)

geerky42 (geerky42):

\[\frac{ \cancel{(a - 4)} }{ a\cancel{(a - 4)} } \times \frac{ 2a }{ (a + 3)(a - 1) }\]\[\dfrac{1}{a}\times\frac{ 2a }{ (a + 3)(a - 1) } \]

geerky42 (geerky42):

You can also cancel a out.

geerky42 (geerky42):

\[\dfrac{1}{\cancel a}\times\frac{ 2\cancel a }{ (a + 3)(a - 1) } \]

geerky42 (geerky42):

Now you are left with \(\dfrac2{(a+3)(a-1)}\)

OpenStudy (anonymous):

okay cool I didn't know that

geerky42 (geerky42):

That's weird, it's not in options.

geerky42 (geerky42):

Or is it?

OpenStudy (anonymous):

I don't know it doesn't show in my text book

geerky42 (geerky42):

What computer are you using?

OpenStudy (anonymous):

HP

geerky42 (geerky42):

Then there should be PrtScr on your keyboard, right? you can take screenshot by press that

geerky42 (geerky42):

then open MS paint and paste

OpenStudy (anonymous):

I know that but I broke the button

geerky42 (geerky42):

ah okay.

geerky42 (geerky42):

well, I don't know what to do from here, sorry.

OpenStudy (anonymous):

it okay thanks for your help